find the x and y intercepts; do not graph.
x-2y=6
and
2x+3y=27

Answer by stanbon(75887)

find the x and y intercepts; do not graph.
x-2y=6
x-intercept:
Let y = 0, then x = 6
--
y-intercept:
Let x = 0, then y = -3
------------------------------
and
2x+3y=27
x-intercept:
Let y = 0, then x = 27/2
---
y-intercept:
Let x = 0, then 27/3 = 9
==============================
Cheers,
Stan H.